OpenStudy (anonymous):

Which expression shows that 11 + 7 = 5 + 13? A. (5 + 6) + 7 = 5 + (6 + 7) B. 6 + 6 + 6 = 18 C. 5 + 9 + 4 = 6 + 7 + 5 D. (11 + 4) + 3 = (5 + 7) + 6

OpenStudy (anonymous):

help

OpenStudy (anonymous):

A. because you can break up the 11 to be 5+6 and the 13 to 6+7 to show that 5+6+7=5+6+7=18

OpenStudy (anonymous):

thanks